Consider two bonds, a 3-year bond paying an annual coupon of 6.30% and a 10-year bond also with an annual coupon of 6.30%. Both currently sell at a face value of $1,000. Now suppose interest rates rise to 9%. a. What is the new price of the 3-year bonds? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Bond price b. What is the new price of the 10-year bonds? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Bond price c. Which bonds are more sensitive to a change in interest rates? Long-term bonds Short-term bonds
Answer #1

a) Coupon = 6.3%*1000 = 63
Par Value =1000
Maturity = 3 years
Price = PV of Coupon + PV of Par Value = 63*(1-(1+9%)-3)/9% + 1000/(1+9%)3 = 931.66

b) Price = PV of Coupon + PV of Par Value = 63*(1-(1+9%)-10)/9% + 1000/(1+9%)10 = 826.72

c)Option a long term bonds

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well